OpenStudy (anonymous):

A two-pen corral is to be built. The outline of the corral forms two identical adjoining rectangles. If there is120 ft of fencing available, what dimensions of the corral will maximize the enclosed area?

OpenStudy (anonymous):

120=2x+2y A=xy A=(60-y)y A=60y-y^2 A'=60-y you set the derivative to 0 so i get a critical number 60 I'm stuck here and don't know where to go because the minimun y can be is 0 and wouldn't the maximum be 60..? but if i plug in 60, i get 0 for the area..... :(